What is the distance between two protons so that their Coulomb potential is barrier potential 200 keV 



1
Expert's answer
2021-06-17T15:16:09-0400

Gives

V=200Kev

"V=\frac{Kz_1z_2e^2}{r}"

For coulamb potential (Alpha partical)


"z_1=z_2=2"

"r=\frac{kz_1z_2e^2}{V}"

Put value


"r=\frac{9\times10^9\times2\times 2\times(1.6\times10^{-19})^2}{200\times1.6\times10^{-16}}=0.288\times10^{-10}"